### v2

# Plug-and-Play Linear Attention with Provable Guarantees for Training-Free Image Restoration

Abstract
[…]
Index Terms—  Attention, Image restoration, Training-free, Plug-and-Play, Nyström approximation.
1 Introduction
Transformers have recently demonstrated strong performance across a range of low-level vision tasks, including image denoising, deblurring [21], super-resolution [8], and dehazing [15]. These models rely on a self-attention mechanism that captures long-range dependencies by computing pairwise similarity between features corresponding to image patches. Let $\mathbf{X}\in\mathbb{R}^{N\times d_{i}}$ be the input feature map corresponding to $N$ tokens (e.g., pixels or patches), where each token is represented by a $d_{i}$-dimensional feature vector. To compute self-attention [18], the input $\mathbf{X}$ is first projected into query, key, and value representations:
$\mathbf{Q}=\mathbf{X}\mathbf{W}_{Q},\quad\mathbf{K}=\mathbf{X}\mathbf{W}_{K},\quad\mathbf{V}=\mathbf{X}\mathbf{W}_{V},$
where $\mathbf{W}_{Q},\mathbf{W}_{K}\in\mathbb{R}^{d_{i}\times d}$, $\mathbf{W}_{V}\in\mathbb{R}^{d_{i}\times d_{v}}$ are learned projection matrices. Then, the attention matrix $\mathbf{A}\in\mathbb{R}^{N\times N}$, its row-normalized form $\mathbf{S}\in\mathbb{R}^{N\times N}$, and the self-attention output $\mathbf{O}\in\mathbb{R}^{N\times d_{v}}$ are computed as:
$\mathbf{A}=\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}},\qquad\mathbf{S}=\operatorname{softmax}(\mathbf{A}),\qquad\mathbf{O}=\mathbf{S}\mathbf{V},$ (1)
where $\operatorname{softmax}(\mathbf{A})$ denotes independently applying softmax to each row of $\mathbf{A}$.
However, the attention operation in (1) incurs a quadratic complexity in the number of input tokens $N$, which corresponds to the number of pixels or patches within a subregion of a frame [18]. In particular, the time and storage complexity for computing (1) is $\mathcal{O}(N^{2}d)$ and $\mathcal{O}(N^{2}+Nd)$, respectively, dominated by storing and computing the large attention matrices $\mathbf{A}$ and $\mathbf{S}$. Currently, most vision models employ Multi-Head Self-Attention (MHSA) [18], an extension of self-attention. In MHSA, the embedding dimension $d$ is divided into $h$ attention heads, each of dimension $d_{h}$, such that $d=hd_{h}$. For each head, self-attention outputs are independently computed, concatenated, and linearly projected to form the MHSA output. The time and storage complexity of MHSA is quadratic in $N$, similar to self-attention. This computational bottleneck makes transformers less suitable for real-time or resource-constrained settings, where convolutional networks are still often preferred due to their computational efficiency [10, 14].
Several approaches approximate self-attention via low-rank [20, 3, 23] or sparse [6] representations to reduce quadratic complexity, but these methods often require retraining. Linformer [20] projects keys/values into lower-dimensional subspaces via learned linear maps, achieving linear complexity in $N$. Performer [3] replaces the softmax kernel with random feature approximation, resulting in linear complexity and unbiased attention estimates. Nyströmformer [23] heuristically reconstructs attention using landmark-based sampling. Axial Attention [7] factorizes 2D attention into 1D along spatial axes, reducing cost from $H^{2}W^{2}$ to $HW(H+W)$, where the number of tokens is $N=HW$. FlashAttention [4] and Triton [17] preserve exact attention while optimizing execution through custom fused GPU kernels, but are hardware-specific and difficult to integrate into pre-trained models.
Interestingly, research on low-rank attention has seen reduced momentum since 2022, with the focus shifting to hardware-aware accelerations. In this work, we revisit classical low-rank approximations that can act as a plug-and-play replacement for attention modules in transformer-based vision models. Specifically, we propose PnP-Nystra, a training-free, plug-and-play approximation of self-attention using the Generalized Nyström approximation [1, 11, 19]. The core contributions of our method are summarized as follows:
1) Linear attention: Unlike MHSA, which incurs quadratic time and memory complexity in the number of tokens $N$, PnP-Nystra achieves linear scaling in both time and memory.
2) Impressive acceleration: PnP-Nystra serves as a drop-in replacement for standard self-attention in existing window-attention transformer backbones for image restoration, enabling substantial inference-time acceleration without modifying the pretrained weights. In our experiments, the GPU inference speedup ranges from $1.8\times$ to $3.6\times$ across restoration tasks, while the gains are even larger on CPU inference.
3) Robust performance without retraining: Despite significant acceleration, the reduction in restoration quality is minimal. For instance, in image denoising, PnP-Nystra achieves more than $3.5\times$ speedup on CPU with <1<1 dB drop in PSNR and <0.03<0.03 drop in SSIM.
To the best of our knowledge, we are the first to show that a purely drop-in linear-attention replacement can remain close to the original pretrained self-attention outputs without any finetuning, while retaining provable approximation guarantees.
2 Proposed Method
In (1), by letting $\mathbf{G}=\exp(\mathbf{A})$ denote the element-wise exponential of the attention matrix $\mathbf{A}$, the self-attention output $\mathbf{O}\in\mathbb{R}^{N\times d_{v}}$ can be reformulated as
$\mathbf{O}=\left(\mathbf{G}\mathbf{V}\right)\oslash\left(\mathbf{G}\mathbf{1}_{N}\right),$ (2)
[…]
$\mathbf{G}_{ij}=\exp\left(\frac{\boldsymbol{q}_{i}^{\top}\boldsymbol{k}_{j}}{\sqrt{d}}\right).$
We propose to leverage the low-rank structure of the kernel matrix $\mathbf{G}$ via a generalized Nyström method adapted for non-symmetric matrices. Originally developed for approximating solutions to functional eigenvalue problems [12, 2], the Nyström method was later extended to efficiently estimate eigenvectors and construct low-rank matrix approximations [16, 22, 5]. To apply the Nyström approximation, we first select $m\ll N$ landmark query and key vectors, denoted by $\{\boldsymbol{\bar{q}}_{i}\}_{i=1}^{m}$ and $\{\boldsymbol{\bar{k}}_{i}\}_{i=1}^{m}$. Let $\bar{\mathbf{Q}}\in\mathbb{R}^{m\times d}$ and $\bar{\mathbf{K}}\in\mathbb{R}^{m\times d}$ denote the matrices formed by stacking these landmark vectors. We define the extended query and key matrices:
$\widetilde{\mathbf{Q}}=\begin{bmatrix}\bar{\mathbf{Q}}\\ \mathbf{Q}\end{bmatrix},\qquad\widetilde{\mathbf{K}}=\begin{bmatrix}\bar{\mathbf{K}}\\ \mathbf{K}\end{bmatrix}$.
We then define the extended kernel matrix $\widetilde{\mathbf{G}}\in\mathbb{R}^{(m+N)\times(m+N)}$ as:
$\widetilde{\mathbf{G}}=\exp\left(\frac{\widetilde{\mathbf{Q}}\widetilde{\mathbf{K}}^{\top}}{\sqrt{d}}\right)=\begin{bmatrix}\mathbf{G}_{A}&\mathbf{G}_{U}\\ \mathbf{G}_{L}&\mathbf{G}\end{bmatrix}$, (3)
where, the core matrix $\mathbf{G}_{A}\in\mathbb{R}^{m\times m}$ captures similarities between landmark queries and keys:
$\mathbf{G}_{A}=\exp\left(\frac{\bar{\mathbf{Q}}\bar{\mathbf{K}}^{\top}}{\sqrt{d}}\right).$ (4)
The cross-similarity matrices are defined as $\mathbf{G}_{L}\in\mathbb{R}^{N\times m}$ (queries with landmark keys) and $\mathbf{G}_{U}\in\mathbb{R}^{m\times N}$ (landmark queries with keys):
$\mathbf{G}_{L}=\exp\left(\frac{\mathbf{Q}\bar{\mathbf{K}}^{\top}}{\sqrt{d}}\right),\quad\mathbf{G}_{U}=\exp\left(\frac{\bar{\mathbf{Q}}\mathbf{K}^{\top}}{\sqrt{d}}\right).$ (5)
Applying Nyström approximation to the extended matrix $\widetilde{\mathbf{G}}$, the original kernel matrix $\mathbf{G}$ is approximated as:
$\widehat{\mathbf{G}}=\mathbf{G}_{L}\mathbf{G}_{A}^{\dagger}\mathbf{G}_{U},$ (6)
where $\mathbf{G}_{A}^{\dagger}$ denotes the Moore-Penrose pseudoinverse. Substituting (6) into (2), the attention output is approximated as:
$\widehat{\mathbf{O}}=\left(\widehat{\mathbf{G}}\mathbf{V}\right)\oslash\left(\widehat{\mathbf{G}}\mathbf{1}_{N}\right).$ (7)
We refer to the approximation in (7) as PnP-Nystra, our method for approximating self-attention. The complete steps for computing PnP-Nystra are summarized in Algorithm 1.
Algorithm 1 PnP-Nystra in (7): Self-Attention Approximation
0: Queries $\mathbf{Q}\in\mathbb{R}^{N\times d}$, Keys $\mathbf{K}\in\mathbb{R}^{N\times d}$, Values $\mathbf{V}\in\mathbb{R}^{N\times d_{v}}$, number of landmarks $m$
(Step 1) Select $m$ landmark query and key vectors to form landmark matrices $\bar{\mathbf{Q}}\in\mathbb{R}^{m\times d},\bar{\mathbf{K}}\in\mathbb{R}^{m\times d}$
(Step 2) Compute core matrix $\mathbf{G}_{A}\in\mathbb{R}^{m\times m}$ in (4)
(Step 3) Compute cross-similarity matrices $\mathbf{G}_{L}\in\mathbb{R}^{N\times m}$ and $\mathbf{G}_{U}\in\mathbb{R}^{m\times N}$ in (5)
(Step 4) Compute pseudoinverse $\mathbf{G}_{A}^{\dagger}\in\mathbb{R}^{m\times m}$
(Step 5) Compute premultiplier $\mathbf{P}\in\mathbb{R}^{N\times m}$ as $\mathbf{G}_{L}\mathbf{G}_{A}^{\dagger}$
(Step 6) Compute numerator $\mathbf{O}_{N}=\mathbf{P}\,(\mathbf{G}_{U}\mathbf{V})\in\mathbb{R}^{N\times d_{v}}$ and compute denominator $\mathbf{O}_{D}=\mathbf{P}\,(\mathbf{G}_{U}\mathbf{1}_{N})\in\mathbb{R}^{N\times 1}$
(Step 7) Normalize $\widehat{\mathbf{O}}=\mathbf{O}_{N}\oslash\mathbf{O}_{D}$ (row-wise division)
return Output $\widehat{\mathbf{O}}\in\mathbb{R}^{N\times d_{v}}$
Note that, while the Nyströmformer [23] also employs the Nyström method, it operates on the softmax-transformed attention matrix $\mathbf{S}$. However, the softmax operation prevents $\mathbf{S}$ from being naturally partitioned into submatrices - a core requirement for Nyström-based approximation. As a result, the authors resort to a heuristic strategy, which makes their approximation devoid of guarantees such as those in Lemma 2.1.
2.1 Approximation error
We analyze the spectral norm error of the Nyström-based kernel approximation used in PnP-Nystra, where the original matrix $\mathbf{G}$ is approximated as $\widehat{\mathbf{G}}=\mathbf{G}_{L}\mathbf{G}_{A}^{\dagger}\mathbf{G}_{U}$.
Lemma 2.1 (Spectral norm error bound).
Assume $\mathbf{G}_{A}$ is nonsingular. Then,
$\|\mathbf{G}-\mathbf{G}_{L}\mathbf{G}_{A}^{\dagger}\mathbf{G}_{U}\|_{2}\leq C\,\frac{\sigma_{m+1}(\widetilde{\mathbf{G}})}{\sigma_{m}(\mathbf{G}_{A})},$
[…]
The proof follows from Lemma 5.10 in [1]. The error bound in Lemma 2.1 indicates that the approximation error decreases when the singular values of $\widetilde{\mathbf{G}}$ decay rapidly. In particular, if the number of landmark points $m$ approaches the rank of $\widetilde{\mathbf{G}}$, then $\sigma_{m+1}(\widetilde{\mathbf{G}})\to 0$, leading to tighter approximations. The following result formalizes this by guaranteeing perfect reconstruction when the number of landmark points equals the rank of the extended kernel matrix.
Corollary 2.2 (Exact recovery).
If $\operatorname{rank}(\widetilde{\mathbf{G}})=m$, then $\sigma_{m+1}(\widetilde{\mathbf{G}})=0$, and the Nyström approximation becomes exact, i.e., $\mathbf{G}=\mathbf{G}_{L}\mathbf{G}_{A}^{\dagger}\mathbf{G}_{U}$.
We next empirically validate the low-rank property of the kernel matrix in pre-trained vision models, which is indeed the assumption underlying Lemma 2.1 and Corollary 2.2. Fig. 1 illustrates the singular value spectrum of the attention matrices from two distinct pre-trained models. In both cases, the rapid singular value decay confirms the low-rank structure of the attention matrices.
(a) SwinIR [8]
(b) Uformer-B [21]
Fig. 1: Variation of the top $50$ singular values of attention maps ($N=32^{2}$) from SwinIR and Uformer, averaged over all heads and layers. The steep decay within the first 2020 singular values highlights the low-rank structure of the attention matrices.
2.2 Computational Details
Standard self-attention has time complexity $\mathcal{O}(N^{2}d)$ and memory $\mathcal{O}(N^{2}+Nd)$. By avoiding explicit $N\times N$ attention, PnP-Nystra reduces the overall complexity to $\mathcal{O}(Nm(d+d_{v}))$ time and $\mathcal{O}(m^{2}+N(m+d_{v}))$ memory for $m\ll N$ (Table 1).
Table 1: Per-step time/space complexity of PnP-Nystra.
Step 2 3 4 5 6 7
Time $\mathcal{O}(m^{2}d)$ $\mathcal{O}(Nmd)$ $\mathcal{O}(m^{3})$ $\mathcal{O}(Nm^{2})$ $\mathcal{O}(Nm(d_{v}{+}1))$ $\mathcal{O}(Nd_{v})$
Space $\mathcal{O}(m^{2})$ $\mathcal{O}(Nm)$ $\mathcal{O}(m^{2})$ $\mathcal{O}(Nm)$ $\mathcal{O}(N(d_{v}{+}1))$ in-place
We form landmarks by average pooling over non-overlapping windows as in [9]. For numerical stability, we replace PyTorch’s pseudoinverse with the iterative method of [13]. Steps 2–3 evaluate exponential kernels, which may overflow and destabilize the Nyström computation. We adopt a softmax-style row-max stabilization by introducing diagonal matrices $\mathbf{M}_{L}\in\mathbb{R}^{N\times N}$ and $\mathbf{M}_{U}\in\mathbb{R}^{m\times m}$ as
$(\mathbf{M}_{L})_{ii}$ $=\exp(m_{L}(i)),$ $\quad m_{L}(i)$ $=\max_{j}\Big(\frac{\mathbf{Q}\bar{\mathbf{K}}^{\top}}{\sqrt{d}}\Big)_{ij}.$
$(\mathbf{M}_{U})_{ii}$ $=\exp(m_{U}(i)),$ $\quad m_{U}(i)$ $=\max_{j}\Big(\frac{\bar{\mathbf{Q}}\mathbf{K}^{\top}}{\sqrt{d}}\Big)_{ij}.$
This rescales each row of $\mathbf{G}_{L}$ and $\mathbf{G}_{U}$ by its maximum log-score. Using $\mathbf{M}_{L}\mathbf{M}_{L}^{-1}=\mathbf{I}$ and $\mathbf{M}_{U}\mathbf{M}_{U}^{-1}=\mathbf{I}$, we rewrite
$\widehat{\mathbf{G}}=\mathbf{M}_{L}\mathbf{M}_{L}^{-1}\mathbf{G}_{L}\,\mathbf{G}_{A}^{\dagger}\,\mathbf{M}_{U}\mathbf{M}_{U}^{-1}\mathbf{G}_{U}=\mathbf{M}_{L}\,\mathbf{G}_{L}^{n}\,(\mathbf{G}_{A}^{n})^{\dagger}\,\mathbf{G}_{U}^{n},$
with normalized blocks
$\mathbf{G}_{L}^{n}:=\mathbf{M}_{L}^{-1}\mathbf{G}_{L},\quad\mathbf{G}_{A}^{n}:=\mathbf{M}_{U}^{-1}\mathbf{G}_{A},\quad\mathbf{G}_{U}^{n}:=\mathbf{M}_{U}^{-1}\mathbf{G}_{U}.$
Equivalently, $\widehat{\mathbf{G}}=\mathbf{M}_{L}\widehat{\mathbf{G}}^{n}$, where $\widehat{\mathbf{G}}^{n}:=\mathbf{G}_{L}^{n}(\mathbf{G}_{A}^{n})^{\dagger}\mathbf{G}_{U}^{n}$ , and the stabilized output in (7) is given as
$\widehat{\mathbf{O}}=(\mathbf{M}_{L}\widehat{\mathbf{G}}^{n}\mathbf{V})\oslash(\mathbf{M}_{L}\widehat{\mathbf{G}}^{n}\mathbf{1}_{N})=(\widehat{\mathbf{G}}^{n}\mathbf{V})\oslash(\widehat{\mathbf{G}}^{n}\mathbf{1}_{N}).$ (8)
In practice, Algorithm 1 computes $\widehat{\mathbf{O}}$ using the stabilized form in (8) instead of (7).
Fig. 2: Single-image dehazing. Top: hazy input and restored outputs. Bottom: ground truth and pixel-wise error maps (scaled to [0–255]). PnP-Nystra matches the pretrained model closely while producing cleaner dehazing than other training-free linear attention baselines, as evidenced by metrics (PSNR(dB), SSIM) and reconstruction of trees and houses in the zoomed area.
3 Experiments
We evaluate drop-in replacement capabilities of PnP-Nystra in existing restoration models, followed by an ablation study of its algorithmic parameters. PSNR and SSIM metrics for the experiments are reported, alongside runtime and speedup on both GPU and CPU platforms. All timings are averaged over multiple inference runs per test dataset. Experiments are conducted using PyTorch $2.7$ with CUDA $12.2$, with GPU and CPU benchmarks performed on NVIDIA RTX $4090$ and Intel Xeon Gold processors, respectively. The implementation is available at: https://anonymous.4open.science/r/PnP_Nystra-0F31.
3.1 PnP-Nystra for restoration
In this section, we evaluate PnP-Nystra as a drop-in replacement for MHSA in three state-of-the-art restoration Transformers: SwinIR [8], Uformer-B [21], and Dehazeformer [15]. We keep all pretrained projection weights $(\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V})$ fixed and perform no finetuning. Unless stated otherwise, we use $m=16$ landmarks and 4466 pseudoinverse iterations, chosen to provide strong performance across tasks. We compare against other training-free linear attention substitutes, such as Performer [3] and Nyströmformer [23]. Our intent is not exhaustive restoration benchmarking, but to quantify the inference-time trade-off of attention approximation within these high-performing pretrained baselines. Speedups over the original attention are summarized in Table 2.
Table 2: Timing comparison per attention block for original model and PnP-Nystra. GPU/CPU time (in ms) is reported.
Resolution Scale Original PnP-Nystra Speedup
SwinIR for Super-Resolution
$88\times 288 22$ $524.61$ / $17916$ $235.76$ / $8139$ $23\times / 2.2$$20\times$
$160\times 160$} 4{4} $159.96$ / $5298$ $65.46$ / $1688$ $2.44\times$ / $3.13\times$
$96\times 96$ 88 $59.18$ / $2274$ $29.11$ / $652$ $2.03\times$ / $3.49\times$
Uformer-B for Denoising
$512\times 512$ $218.01$ / $6958$ $67.26$ / $1853$ $3.24\times$ / $3.75\times$
$256\times 256$ $81.81$ / $2610$ $22.64$ / $372$ $3.61\times$ / $7.05\times$
Uformer-B for Deblurring
$768\times 768$} $285.46$ / $9098$ $155.19$ / $4942$ $\times / 1.84×1.84\times$
Dehazeformer for Dehazing
$416\times 576$ $143.92$ / $5106$ $51.71$ / $1689$ $2.78\times$ / $3.02\times$
$576\times 576$ $183.53$ / $6938$ $70.05$ / $2558$ $2.62\times$ / $2.71\times$
$736\times 576$ $261.80$ / $10192$ $105.02$ / $4157$ $2.49\times$ / $2.45\times$
3.1.1 Dehazing
We evaluate Dehazeformer [15] with PnP-Nystra as a drop-in replacement for MHSA on single-image dehazing. We use a window size of $32$ (i.e., $N=32^{2}$ tokens) and run $5$ pseudoinverse iterations for all dehazing experiments. Table 3 reports PSNR/SSIM on different datasets, where PnP-Nystra consistently outperforms other training-free linear attention substitutes, producing accuracy close to (and in some cases exceeding) the original pretrained Dehazeformer. This is achieved at a $2.5\times$ to $2.8\times$ speedup as shown in the Table. 2. Qualitative results in Fig. 2 show that PnP-Nystra preserves fine details and contrast on par with the original model, while producing cleaner haze removal than other training-free approximations.
Table 3: Accuracy Comparison for Image Dehazing (PSNR / SSIM) with other linear attention methods
Dataset Original Performer Nystromformer PnP-Nystra
Reside-Out $26.51$ / $0.9512$ $25.26$ / $0.9328$ $23.57$ / $0.9266$ 26.34 / 0.9510
RSHaze $32.80$ / $0.9261$ $29.00$ / $0.8957$ $33.18$ / $0.9363$ 33.58 / 0.9405
RESIDE-6K $24.79$ / $0.9352$ $21.00$ / $0.8834$ $21.37$ / $0.8962$ 23.08 / 0.9163
3.1.2 Denoising and Deblurring
[…]
Table 4: Accuracy comparison (PSNR / SSIM) for Uformer-B with training-free linear attention substitutes.
Dataset Original Performer Nyströmformer PnP-Nystra
Denoising
SIDD $38.89$ / $0.8950$ $34.80$ / $0.8713$ $37.46$ / $0.8838$ $\mathbf{37.88}$ / $\mathbf{0.8842}$
BSDS200 $28.05$ / $0.8067$ $24.80$ / $0.7469$ $26.77$ / $0.7829$ $\mathbf{27.17}$ / $\mathbf{0.7838}$
Deblurring
RealBlur-R $33.98$ / $0.9404$ $33.97$ / $0.9384$ $34.01$ / $0.9407$ $\mathbf{34.02}$ / $\mathbf{0.9407}$
3.1.3 Super-resolution
Table 5 compares SwinIR and its PnP-Nystra variant on Set5 and BSDS100 across different scales for a window size of $32$, i.e., $N=32^{2}$. As shown in the Table. 2, PnP-Nystra consistently achieves $2$-$3\times$ speedup on both GPU and CPU, with better runtime benefits at lower resolutions. Even with the speedup, the accuracy drop, measured in PSNR and SSIM, is lowest among the compared linear attention methods in most evaluations in the Table. 5 across different scaling factors.
Table 5: SwinIR super-resolution accuracy (PSNR(dB),SSIM) comparison across scaling factors.
Dataset Scale Original Performer Nyströmformer PnP-Nystra
Set5 22 $36.13$ / $0.9455$ $33.91$ / $0.9322$ $35.15$ / $0.9413$ 35.25 / 0.9416
44 $30.86$ / $0.8729$ $28.11$ / $0.8142$ 29.26 / 0.8467 29.10{29.10} / 0.8437{0.8437}
88 $25.62$ / $0.7439$ $23.53$ / $0.6434$ $24.26$ / $0.6787$ 24.41 / 0.6853
BSDS100 22 $30.98$ / $0.8922$ $29.82$ / $0.8740$ $30.64$ / $0.8821$ 30.71 / 0.8830
44 $26.43$ / $0.7183$ $24.98$ / $0.6815$ $25.81$ / $0.6992$ 25.92 / 0.7008
88 $23.61$ / $0.5680$ $22.73$ / $0.5291$ $23.21$ / $0.5487$ 23.29 / 0.5503
3.2 Ablation Study of PnP-Nystra
In this section, we present an ablation study of PnP-Nystra, with emphasis on the internal hyperparameters.
Scaling of Inference Time with $N$: Fig. 3 presents the CPU inference runtime of the original MHSA module versus PnP-Nystra (averaged over repeated runs), plotted in both linear and logarithmic scales with respect to the token count $N$. In line with the complexity analysis, PnP-Nystra exhibits linear scaling with an increasing number of tokens.
(a) Linear scale
(b) Logarithmic scale
Fig. 3: Inference time vs. token count NN: Unlike MHSA which grows quadratically with NN, PnP-Nystra scales linearly.
Landmarks and Pseudoinverse iterations: We evaluate the effect of varying the number of landmarks $m$ and the number of iterations used for pseudoinverse computation in Step $4$ of Algorithm 1. Table 6 reports PSNR and SSIM drop for the PnP-Nystra variant of Uformer-B for denoising on the SIDD dataset. As expected from Nyström approximation, the approximation performance tends to the original with an increase in the number of landmarks. Additionally, a better estimate of the pseudoinverse leads to improved approximation.
Table 6: Impact of Number of Landmarks and Pseudoinverse Iterations on Approximation Accuracy
Varying # Landmarks (at 6 Iterations) Varying Iterations (at 16 Landmarks)
# Landmarks PSNR Drop SSIM Drop Iterations PSNR Drop SSIM Drop
88 $1.54$ $0.0127$ 11 $2.84$ $0.0167$
$16$ $1.01$ $0.0108$ 33 $1.94$ $0.0138$
$32$ $0.86$ $0.0102$ 66 $1.01$ $0.0108$
Attention map comparison: PnP-Nystra approximates the kernel matrix $\mathbf{G}$ and therefore induces an approximate attention map $\mathbf{S}$. As shown in Fig. 4, the resulting attention distribution closely follows that of the original model, highlighting the same prominent edges and structural regions while maintaining similar suppression over less informative areas.
Fig. 4: Attention map visualization for Uformer-B: (a) original model and (b) PnP-Nystra. Both maps have the same salient structures with strong responses along object boundaries
4 Conclusion and Future Work
We presented PnP-Nystra, a plug-and-play Nyström approximation for self-attention that enables efficient inference in window-based transformer models for image restoration. Experiments show that PnP-Nystra acts as a fast, training-free replacement for self-attention in pre-trained models, offering a viable solution for resource-constrained deployment. Future work will focus on extending PnP-Nystra to global attention in vision transformers and diffusion transformers, and spatiotemporal attention for video restoration.